Prove that the conic passing through the points of intersection of two rectangular hyperbolas is also a rectangular hyperbola.


1
Expert's answer
2022-02-04T08:24:16-0500

Lets say S = 0 and S' = 0 are the two rectangular hyperbolas,

we then have a + b = 0, and a' + b' = 0.

Hence, in the conic

"S-\gamma S'=0"

The sum of the coefficients of x2 and y2

"=(a-\gamma a')+(b-\gamma b')\\=(a+b)+\gamma (a'+b')=0"

Hence the conic

"S-\gamma S'=0"

itself a rectangular hyperbola.